Machinery purchased for $69,600 by Tamarisk Co. in 2016 was originally estimated to have a life of 8 years with a salvage value of $4,640 at the end of that time. Depreciation has been entered for 5 years on this basis. In 2021, it is determined that the total estimated life should be 10 years with a salvage value of $5,220 at the end of that time. Assume straight-line depreciation.

Prepare the entry to correct the prior years' depreciation, if necessary. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Prepare the entry to record depreciation for 2021. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Machinery related account.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1
Prepare the entry to correct the prior years' depreciation, if necessary. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

No Journal entry 0
Prepare the entry to record depreciation for 2021. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Machinery related account.)

Account Titles and Explanation

Debit

Credit

Depreciation expense 4756
Accumulated depreciation-Machinery 4756

Accumulated depreciation = ( 69600-4640/8*5) = 40600

Revised depreciation = (69600-40600-5220)/5 = 4756